# Utilizing Contextual Clues and Role Correlations for Enhancing Document-level Event Argument Extraction

### v2

@@ Abstract
Document-level event argument extraction (EAE) is a vital but challenging task in the domain of information extraction. Most existing approaches focus on the interaction between arguments and event triggers, ignoring two critical points: the information of contextual clues and the semantic correlations among argument roles. In this paper, we propose the CARLG (Contextual Clues Aggregation and Role-based Latent Information Guidance) model, which consists of two modules: Contextual Clues Aggregation (CCA) and Role-based Latent Information Guidance (RLIG), effectively leveraging contextual clues and role correlations for improving document-level EAE. The CCA module adaptively captures and integrates contextual clues by utilizing context attention weights from a pre-trained encoder. The RLIG module captures semantic correlations through role-interactive encoding and provides valuable information guidance with latent role representation. Notably, our CCA and RLIG modules are compact, transplantable and efficient, which introduce no more than 1% new parameters and can be easily equipped on other span-base methods with significant performance boost. Extensive experiments on the RAMS, WikiEvents, and MLEE datasets demonstrate the superiority of the proposed CARLG model. It outperforms previous state-of-the-art approaches by 1.26 F1, 1.22 F1, and 1.98 F1, respectively, while reducing the inference time by 31%. Furthermore, we provide detailed experimental analyses based on the performance gains and illustrate the interpretability of our model.
@@
Event Argument Extraction (EAE), Information Extraction (IE), Natural Language Processing (NLP), Natural Language Understanding (NLU).
journal: Knowledge-Based Systems
@@ 1 Introduction
Within the domain of Information Extraction (IE), document-level event argument extraction (EAE) is a vital task focused on identifying event arguments and their corresponding roles within textual content associated with specific events. This task is a key process of event extraction (EE) xu2023multi ; lu2022explainable ; lv2022trigger and has broad implications across various fields. For instance, it is instrumental in the analysis of social media networks mitchell1974social ; DING2023110687 ; YANG2022109939 , enhancing recommendation systems  lietal2020gaia ; LI2024111124 ; DONG2022108954 , and improving dialogue systems  Zhang2020.08.03.20167569 ; ZHAO2023110927 . As illustrated in Fig. 1, with the event trigger “homicide” of event type Violence, the goal of EAE system is to identify argument “Freddie Gray” as the role victim, “Baltimore” as the place, and “Officer Caesar Goodson Jr.” as the killer.
@@
Compared with sentence-level EAE liuetal2018jointly ; waddenetal2019entity ; tongetal2020improving , document-level EAE faces more complex issues, notably long-range dependencies and the need for inferences across multiple sentences, discussed in research ebner2020multi and lietal2021document . To solve these problems, some prompt-based methods ma2022prompt ; hsu2023ampere design suitable templates for long text, which solves the long-distance dependency problem and extract arguments in a generative slot-filling manner. Additionally, some works utilize graph structures employing heuristic approaches zhang2020two ; pouranbenveysehetal-2022-document , syntactical frameworks xuetal2021document ; liu2022document ; 9204800 , and Graph Neural Networks kipfsemi for the facilitation of reasoning across sentences. However, current mainstream approaches ignore two essential aspects: (a) the contextual clue information; (b) the semantic correlations among roles.
Contextual clues refer to text elements in the document, excluding the target arguments themselves, which provide beneficial guidance for predicting complex argument roles. These clues play a significant role in event argument extraction by offering valuable information. For instance, in Figure 1, the contextual clues, such as death, accused, and trial, serve as important indicators for identifying the arguments “Freddie Gray” and “Officer Caesar Goodson Jr” within the event Violence. By considering these contextual clues, the model can leverage wider context information to make more informed and accurate predictions. However, many prior studies xu2022two ; lietal2021document ; liu-etal-2023-enhancing-document just rely on the pre-trained transformer encoder to capture global context information in text implicitly, ignoring the necessity of context clues information that is highly relevant to the entity zhou2021document and target event ebner2020multi . Therefore, in this paper, we propose a Contextual Clues Aggregation (CCA) module, which leverages the contextual attention product from the pre-trained transformer encoder to aggregate the event-specific contextual clues for each argument, enhancing the candidate argument representation by incorporating broader and relevant contextual information.
@@
Figure 1: A document from RAMS ebner2020multi dataset. The event violence has the trigger of Homicide and there are three distinct arguments, each assuming different roles. We use the color red to emphasize specific words, representing contextual clues that are significant for the extraction of arguments.
Within the context of real-world document-level EAE task, numerous roles exhibit semantic correlations. For example, as illustrated in Figure 1, there exists a strong semantic correlation between the roles of killer and victim, serving as beneficial guidance for extracting arguments associated with these roles within the target event Violence. In addition, many roles co-occur in multiple events, indicating strong potential for close semantic correlations among them. For a more in-depth analysis, we perform a frequency count and visualize the co-occurrence of the top 15 roles in the RAMS dataset, as shown in Figure 2. Notably, roles such as attacker, target and instrument show high co-occurrence frequency, suggesting their strong semantic correlations compared with other roles. Therefore, in this paper, we propose a Role-based Latent Information Guidance (RLIG) module, comprising role-interactive encoding, the fusion of role information and latent role guidance. Specifically, in role-interactive encoding, we add roles into the input sequence, enabling role embeddings to learn both semantic information and latent correlations among roles through self-attention mechanism. Subsequently, the latent role embeddings are combined with candidate argument spans via pooling and concatenation operations. Finally, we achieve alignment of argument spans, triggers, and roles within a unified space through a novel alignment mechanism utilizing Tucker Decomposition, tucker1964extension , which provides valuable information guidance for document-level EAE.
In summary, to better integrate and leverage our two techniques, we improve existing span-based baselines zhang2020two ; DBLP:journals/corr/abs-1904-05255 for document-level EAE, which has small parameter size and high efficiency. Then the CARLG (Contextual Clues Aggregation and Role-based Latent Information Guidance) model is proposed, by equipping CCA and RLIG modules on our baseline for the aforementioned two crucial points respectively. Significantly, each module leverages the fine-organized attentive weights obtained from the transformer-based pre-trained model, introducing less than 1% of new parameters. Additionally, these two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks (as shown in our Sec. 5.2), which are compact and transplantable. The computational cost of our model is also lower compared with the recent state-of-the-art approach TSAR, saving no less than 31% of inference time. ur contributions are summarized as follows:
  • 1.
    An enhanced span-based baseline for document-level EAE tasks is introduced, which demonstrates both parameter and computation efficiency, significantly outperforming existing span-based baselines
  • 2.
    We propose a Contextual Clues Aggregation (CCA) module, which adaptively identifies and aggregates the contextual clue information and enhances the representation of candidate arguments by incorporating event-specific contextual information.
  • 3.
    We introduce a Role-based Latent Information Guidance (RLIG) module designed to offer guidance based on latent role information, capturing the semantic correlations among roles.
  • 4.
    Extensive experiments on the public RAMS and WikiEvents lietal2021document datasets demonstrate that the proposed approach combined effectively with various span-based document-level EAE models and lead to significant performance improvements. Our CARLG consistently improves EAE performance on the biological dataset MLEE 10.1093/bioinformatics/bts407 , showing the generality of our approach in different domains.
@@ 2.1 Event Argument Extraction
Event argument extraction (EAE) is pivotal for improving event semantic understanding by identifying event arguments within a context, which constitutes a fundamental step in various downstream applications, including question answering liu2022document ; 6544211 ; 9257099 , text summarization allahyari2017text ; 8949697 , and assisted decision-making zhang2020effect ; 9339934 .
@@
Previous approaches primarily concentrate on identifying the event trigger and its associated arguments from individual sentences. In chenetal2015event , a neural pipeline model was first proposed for event extraction, which was later extended by nguyengrishman2015event ; liuetal2017exploiting to incorporate recurrent neural networks (RNNs) and convolutional neural networks (CNNs) into the pipeline model.  liuetal2018jointly ; yanetal2019event utilize dependency trees to construct semantic and syntactic dependencies and relations, which captures the dependency of words in a sentence. waddenetal2019entity explores all possible spans and utilizes graph neural networks to construct span graphs for information propagation. Additionally, transformer-based pre-trained models  waddenetal2019entity ; tongetal2020improving ; liuetal2022dynamic achieve impressive performance in event extraction.
@@
However, in real-world situations, elements of an event are often expressed across multiple sentences, leading to the emergence of document-level event extraction (DEE) 6910247 ; 7347400 ; SU2024111180 ; SCABORO2023110675 . DEE seeks to identify event arguments throughout whole documents, addressing the issue of extended dependencieswangetal2022query ; xu2022two . We summarize these recent DEE works into 3 categories: (1) Tagging-based methods. Works like wangetal2021cleve ; ducardie2020document apply the BiLSTM-CRF framework for document-level event extraction (DEE). Zheng zhengetal2019doc2edag introduces a mixed graph and a tracking module to capture the interrelations between events. (2) span-based methods. yangetal2021document proposes an encoder-decoder framework for parallel extraction of structured events. Works ebner2020multi ; zhang2020two predict argument roles for selected text segments under a maximum length restriction. Furthermore, Xu xu2022two suggests a dual-stream encoder complemented by an AMR-guided graph to manage extended dependencies. (3) generation-based methods. lietal2021document approaches the issue through conditional generation, whereas duetal2021grit views it as a sequence-to-sequence endeavor. Additionally, Wei weietal2021trigger redefines the task in terms of a reading comprehension challenge. Furthermore, Ren renetal2022clio integrates argument roles into document encoding to account for multiple role details for nested arguments. ma2022prompt ; liu2022document ; XU2023110375 propose a prompt-based method that utilizes slotted prompts for generative slot-filling to extract arguments.
@@ 2.2 Contextual Clues
Contextual clues refer to the information and cues present in the surrounding context of a given entity, event, or situation clark2016contextual ; zhang2020contextual ; liu2022document . These clues provide additional context and can help in understanding the meaning, significance, or relationships associated with the target entity or event peters2018deep . NC-DRE zhang2022nc introduces an encoder-decoder structure, leveraging a pre-trained encoding model and a graph neural network (GNN) for decoding. This setup facilitates the use of non-entity clues in document-level relation extraction. Concurrently, a method for localized context pooling zhou2021document has been put forward, allowing for the direct application of attention mechanisms from pre-trained language models. This method efficiently leverages the contextual information tied to particular entity pairs, aiding in capturing the beneficial context clues for determining relations. Xu xu2021discriminative uses the self-attention mechanism vaswani2017attention to learn a document-level context representation (DLCRep) which contributes to the improvement of information extraction.
Therefore, contextual clues contain important semantic information, which can provide information guidance for deep mining of semantic associations between entities. In this paper, we are the first to incorporate contextual clues techniques into the event argument extraction task and propose the Contextual Clues Aggregation (CCA) module to explicitly capture contextual clues.
@@ 2.3 Relation Embedding
Many methods in the field of information extraction leverage relation embeddings to enhance the effectiveness of various tasks 9086164 ; 9694521 ; 9123572 . CRE chen-badlani-2020-relation presents a novel approach to relation extraction by utilizing sentence representations as relation embeddings, which are assessed through entity embeddings trained concurrently with the knowledge base. xuetal2022emrel constructs relation representation explicitly and interacts it with entities and context, which achieves excellent results in multi-triple relation extraction.  han2022documentlevel employs relation embeddings as a medium and introduces two sub-tasks focused on co-occurrence prediction, aiming to enhance the retrieval of relational facts.
@@
3 Enhanced Span-based Baseline
To better integrate our two techniques, we first propose an enhanced span-based baseline for document-level EAE, which exhibits small parameter size and high efficiency. We clarify the preliminaries of the document-level EAE task as follows for better understanding our method.
3.1 Preliminaries
We follow the definition of event argument extraction in RAMSv1.0 benchmark ebner2020multi . To better understand this task, we first clarify some important notions.
  • 1.
@@ 3.1 Preliminaries
  • 2.
    Event trigger (trigger): the key term that most explicitly represents the manifestation of the event, e.g., the word homocide in Figure 1.
  • 3.
@@
Figure 3: The architecture or our enhanced baseline.

### v4

# mode = title

@@ Abstract
Document-level event argument extraction is a crucial yet challenging task within the field of information extraction. Current mainstream approaches primarily focus on the information interaction between event triggers and their arguments, facing two limitations: insufficient context interaction and the ignorance of event correlations. Here, we introduce a novel framework named CARLG (Contextual Aggregation of Clues and Role-based Latent Guidance), comprising two innovative components: the Contextual Clues Aggregation (CCA) and the Role-based Latent Information Guidance (RLIG). The CCA module leverages the attention weights derived from a pre-trained encoder to adaptively assimilate broader contextual information, while the RLIG module aims to capture the semantic correlations among event roles. We then instantiate the CARLG framework into two variants based on two types of mainstream EAE approaches. Notably, our CARLG framework introduces less than 1% new parameters yet significantly improving the performance. Comprehensive experiments across the RAMS, WikiEvents, and MLEE datasets confirm the superiority of CARLG, showing significant superiority in terms of both performance and inference speed compared to major benchmarks. Further analyses demonstrate the effectiveness of the proposed modules.
@@
Event Extraction, Information Extraction, Feature Engineering, Natural Language Processing, Knowledge Graph.
@@ 1 Introduction
Within the domain of Information Extraction, document-level event argument extraction (EAE) is a vital task focused on identifying event arguments and their corresponding roles within textual content associated with specific events. This task is a key process of event extraction (Xu et al., 2023b; Lu et al., 2022; Lv et al., 2022) and has broad applications across various fields. For instance, it is instrumental in the analysis of social media networks (Mitchell, 1974; Ding et al., 2023; Yang et al., 2022), enhancing recommendation systems  (Li et al., 2020, 2024a; Dong et al., 2022), and improving dialogue systems  (Zhang et al., 2020a; Zhao et al., 2023). As illustrated in Fig. 1, with the event trigger “homicide” of event type Violence, the goal of EAE system is to identify argument “Freddie Gray” as the role victim, “Baltimore” as the place, and “Officer Caesar Goodson Jr.” as the killer.
@@
Compared with sentence-level EAE (Liu et al., 2018; Wadden et al., 2019; Tong et al., 2020), document-level EAE faces more complex issues, notably long-range dependencies and the need for inferences across multiple sentences, as discussed in research (Ebner et al., 2020) and (Li et al., 2021). To solve these problems, some prompt-based methods (Ma et al., 2022a; Hsu et al., 2023) design suitable templates for long text, which solves the long-distance dependency problem and extract arguments in a generative slot-filling manner. Additionally, some works utilize graph structures employing heuristic approaches (Zhang et al., 2020d; Pouran Ben Veyseh et al., 2022), syntactical frameworks (Xu et al., 2021a; Liu et al., 2022a; Wu et al., 2022), and Graph Neural Networks (Kipf & Welling, ) for the facilitation of reasoning across sentences. However, current mainstream approaches ignore two essential aspects: (a) the contextual clue information; (b) the semantic correlations among roles.
@@
Figure 1: A document from RAMS (Ebner et al., 2020) dataset. The event violence has the trigger of Homicide and there are three distinct arguments, each assuming different roles. We use the color red to emphasize specific words, representing contextual clues that are significant for the extraction of arguments.
Contextual clues refer to text elements in the document, excluding the target arguments themselves, which provide beneficial guidance for predicting complex argument roles. These clues play a significant role in event argument extraction by offering valuable information. For instance, in Figure 1, the contextual clues, such as death, accused, and trial, serve as important indicators for identifying the arguments “Freddie Gray” and “Officer Caesar Goodson Jr” within the event Violence. By considering these contextual clues, the model can leverage wider context information to make more informed and accurate predictions. However, many prior studies (Xu et al., 2022b; Li et al., 2021; Liu et al., 2023b) just rely on the pre-trained transformer encoder to capture global context information in text implicitly, ignoring the necessity of context clues information that is highly relevant to the entity (Zhou et al., 2021) and target event (Ebner et al., 2020). Therefore, in this paper, we propose a Contextual Clues Aggregation (CCA) module, which leverages the contextual attention product from the pre-trained transformer encoder to aggregate the event-specific contextual clues for each argument, enhancing the candidate argument representation by incorporating broader and relevant contextual information.
@@
For real-world document-level EAE task, numerous event roles exhibit semantic correlations. For example, as illustrated in Figure 1, there exists a strong semantic correlation between the roles of killer and victim, serving as beneficial guidance for extracting arguments associated with these roles within the target event Violence. In addition, many roles co-occur in multiple events, indicating strong potential for close semantic correlations among them. For a more in-depth analysis, we perform a frequency count and visualize the co-occurrence of the top 15 roles in the RAMS dataset, as shown in Figure 2. Notably, roles such as attacker, target and instrument show high co-occurrence frequency, suggesting their strong semantic correlations compared with other roles. Therefore, in this paper, we propose a Role-based Latent Information Guidance (RLIG) module, comprising role-interactive encoding, the fusion of role information and latent role guidance. Specifically, in role-interactive encoding, we add roles into the input sequence, enabling role embeddings to learn both semantic information and latent correlations among roles through self-attention mechanism. Subsequently, the latent role embeddings are combined with candidate argument spans via pooling and concatenation operations. Finally, we achieve alignment of argument spans, triggers, and roles within a unified space through a novel alignment mechanism utilizing Tucker Decomposition (Tucker et al., 1964), which provides valuable information guidance for document-level EAE.
In summary, in this paper, we propose the CARLG framework (Contextual Clues Aggregation and Role-based Latent Information Guidance), consisting of CCA module and RLIG module. Each module leverages the fine-organized attentive weights obtained from the transformer-based pre-trained model, introducing less than 1% of new parameters. We instantiate the CARLG framework into two variants: the span-based $\text{CARLG}_{\text{span}}$ and the prompt-based $\text{CARLG}_{\text{prompt}}$, based on two types of current mainstream EAE approaches. Both variants demonstrate significant superiority over leading benchmarks, in terms of both performance and inference speed. Our contributions are summarized as follows:
  • To capture the contextual clues within the text, we propose a Contextual Clues Aggregation (CCA) module. It adaptively identifies and aggregates the contextual clue information and enhances the representation of candidate arguments, achieving sufficient context interaction for document-level EAE.
  • We propose a Role-based Latent Information Guidance (RLIG) module to capture event role correlations, which employs learnable latent role embeddings to enhance the semantic interaction among event roles.
  • We instantiate the CARLG framework into two variants: the $\text{CARLG}_{\text{span}}$ and the $\text{CARLG}_{\text{prompt}}$, based on two types of current mainstream EAE approaches, which shows the adaptability and transferability of our framework.
  • Extensive experiments on the public RAMS and WikiEvents (Li et al., 2021) and MLEE (Pyysalo et al., 2012) datasets demonstrate that both CARLG variants show significant superiority in both performance and efficiency. Comprehensive ablation studies and analyses illustrate the effectiveness of our proposed framework.
@@ 2.1 Event Argument Extraction
Event argument extraction (EAE) is pivotal for improving event semantic understanding by identifying event arguments within a context, which constitutes a fundamental step in various downstream applications, including question answering (Liu et al., 2022a; Sun et al., 2013; Liu et al., 2022c), text summarization (Allahyari et al., 2017; Huang et al., 2021), and assisted decision-making (Zhang et al., 2020b; Sakai & Iiduka, 2022).
@@
Previous approaches primarily concentrate on identifying the event trigger and its associated arguments from individual sentences. In (Chen et al., 2015), a neural pipeline model was first proposed for event extraction, which was later extended by (Nguyen & Grishman, 2015; Liu et al., 2017) to incorporate recurrent neural networks (RNNs) and convolutional neural networks (CNNs) into the pipeline model.  (Liu et al., 2018; Yan et al., 2019) utilize dependency trees to construct semantic and syntactic dependencies and relations, which captures the dependency of words in a sentence. (Wadden et al., 2019) explores all possible spans and utilizes graph neural networks to construct span graphs for information propagation. Additionally, transformer-based pre-trained models  (Wadden et al., 2019; Tong et al., 2020; Liu et al., 2022b) achieve impressive performance in event extraction.
@@
However, in real-world situations, elements of an event are often expressed across multiple sentences, leading to the emergence of document-level event extraction (DEE) (Cai & Chen, 2015; Jiang et al., 2023; Su et al., 2024; Scaboro et al., 2023). DEE seeks to identify event arguments throughout whole documents, addressing the issue of extended dependencies(Wang et al., 2022; Xu et al., 2022b). We summarize these recent DEE works into 3 categories: (1) Tagging-based methods. Works like (Wang et al., 2021; Du & Cardie, 2020a) apply the BiLSTM-CRF framework for document-level event extraction. Zheng (Zheng et al., 2019) introduces a mixed graph and a tracking module to capture the interrelations between events.(Li et al., 2021) approaches the issue through conditional generation, whereas (Du et al., 2021) views it as a sequence-to-sequence endeavor. Additionally, Wei (Wei et al., 2021) redefines the task in terms of a reading comprehension challenge. (2) Span-based methods. (Yang et al., 2021) proposes an encoder-decoder framework for parallel extraction of structured events. Works (Ebner et al., 2020; Zhang et al., 2020d) predict argument roles for selected text segments under a maximum length restriction. Furthermore, Xu (Xu et al., 2022b) suggests a dual-stream encoder complemented by an AMR-guided graph to manage extended dependencies. (3) Prompt-based methods. Ren (Ren et al., 2022) integrates argument roles into document encoding to account for multiple role details for nested arguments. (Ma et al., 2022a; Liu et al., 2022a; Xu et al., 2023c) propose a prompt-based method that utilizes slotted prompts for generative slot-filling to extract arguments. Among them, span-based and prompt-based methods are current mainstream methods, which have been demonstrated superior generalizability and competitive performance (Hsu et al., 2023).
@@ 2.2 Contextual Clues
Contextual clues refer to the information and cues present in the surrounding context of a given entity, event, or situation (Jiang et al., 2022; Zhang et al., 2020c; Liu et al., 2022a; Li et al., 2024b). These clues provide additional context and can help in understanding the meaning, significance, or relationships associated with the target entity or event (Peters et al., 2018). NC-DRE (Zhang & Cheng, 2022) introduces an encoder-decoder structure, leveraging a pre-trained encoding model and a graph neural network (GNN) for decoding. This setup facilitates the use of non-entity clues in document-level relation extraction. Concurrently, a method for localized context pooling (Zhou et al., 2021) has been put forward, allowing for the direct application of attention mechanisms from pre-trained language models. This method efficiently leverages the contextual information tied to particular entity pairs, aiding in capturing the beneficial context clues for determining relations. Xu (Xu et al., 2021b) uses the self-attention mechanism (Vaswani et al., 2017) to learn a document-level context representation which contributes to the improvement of information extraction.
Therefore, contextual clues contain important semantic information, which can provide information guidance for deep mining of semantic associations between entities. In this paper, we are the first to incorporate contextual clues techniques into the event argument extraction task and propose the Contextual Clues Aggregation module to explicitly capture contextual clues.
@@ 2.3 Relation Embedding
Many methods in the field of information extraction leverage relation embeddings to enhance the effectiveness of various tasks (Ji et al., 2022; Sun et al., 2023; Lu & Chow, 2022). CRE (Chen & Badlani, 2020) presents a novel approach to relation extraction by utilizing sentence representations as relation embeddings, which are assessed through entity embeddings trained concurrently with the knowledge base. (Xu et al., 2022a) constructs relation representation explicitly and interacts it with entities and context, which achieves excellent results in multi-triple relation extraction.  (Han et al., 2022) employs relation embeddings as a medium and introduces two sub-tasks focused on co-occurrence prediction, aiming to enhance the retrieval of relational facts.
@@
Figure 3: Contextual Clues Aggregation Module. The core process of the Role-based Latent Information Guidance algorithm is also similar to this.
3 Preliminaries
3.1 Important Notions
We follow the definition of event argument extraction in RAMSv1.0 benchmark (Ebner et al., 2020). To better understand this task, we first clarify some important notions.
@@ 3.1 Important Notions
  • Event trigger (trigger): the key term that most explicitly represents the manifestation of the event, e.g., the word homicide in Figure 1.
@@
3.2 Task Formulation